How quickly can you do this? Fill appropriate sign. ( ‘’) 
 6 / 10 ☐ 3 / 5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

6 / 10 = (3 × 2) / (5 × 2)
=3 / 5
∴6 / 10 = 3 / 5

Thanks